The length of a rec is 3 more than twice the width. What are the dimensions if the perimeter is 51
Answer by oberobic(2304)

L = 2W+3
.
2(L+W) = 51
.
substitute
.
2(2W+3+W) = 51
4W +6 +2W = 51
6W = 45
W = 45/6 = 7 1/2
L = 2W+3 = 18
P = 25 1/2 *2 = 51
.
Answer: L = 18, W = 7 1/2
.
Done.